# Strategy for measuring energy spectra of air shower components using a hybrid detector station

### v1

@@ 2.1 Simulation Framework
In this work, the extensive air showers have been simulated using CORSIKA v7.7410 [21], and the detector response is obtained using a stand-alone program based on the Geant4 toolkit (v10.5.1) [22, 23]. In general, a set of 5000 CORSIKA proton-induced showers with a primary energy of E eV and a zenith angle of was used. A thinning algorithm with was applied and hadronic interactions were handled by FLUKA and EPOS-LHC at low and high energies, respectively. In each event, the shower core is injected at a fixed position relative to a specific detector station.
Dedicated datasets were generated to assess the impact of key shower parameters (energy, zenith angle, and core position). For energy resolution, 5000 showers were simulated with energies ranging from eV to eV. For angular resolution, 2000 showers were simulated with a zenith angle of .The shower core position was randomly distributed within a square for each event to investigate the impact of core position reconstruction.
@@ 2.2 Method to evaluate the sensitivity to the particle energy spectrum
In this work, we adopt a phenomenological approach to modify the energy spectra of the electromagnetic and muonic components at ground level, allowing us to assess the detectors’ sensitivity to energy variations. The energy distribution of particles as collected in the multi-hybrid station is adjusted to alter the slopes of the low- or high-energy power-law tails while preserving the event multiplicity.
The original particle energy distribution is obtained by processing CORSIKA-generated air showers down to the detector level using the aforementioned Geant4-based framework. Figure 3 illustrates examples of these distributions for the electromagnetic and muonic shower components. To modify the distributions, a function of the following form is applied to the original spectrum:
This function modifies the spectrum’s tail in a controlled manner. The parameters and define the histogram bins corresponding to the region of the spectrum to be modified while controls the intensity of the modification and runs over the bins being modified. The choice of a monotonically increasing function for weighting the original spectrum is deliberate. Applying a flat increase within a limited region would introduce abrupt spikes at the transition between the modified and unmodified regions. This functional form ensures a smooth transition between the modified and unmodified regions, preserving the spectrum’s overall shape and features.
@@ 3.2 Low-energy tail of the muonic component
Following the approach used for modifying the high-energy tail of the electromagnetic particle energy distribution (Section 3.1), we considered a range of modifications with magnitudes from to . For each case, the estimators and were calculated for both SSD-WCD and WCD-RPC difference plots. The results are presented in Figure 10.
For both difference plots, the values of show clear separation across the different modification intensities, validating the applicability of this method for this configuration. However, for the SSD-WCD difference plot, the values of are less consistent, particularly at lower values of . This is likely due to reduced statistical precision, a consequence of positioning the station farther from the shower core. Nevertheless, for the WCD-RPC difference plots, the values of remain consistent across all modification intensities, emphasizing the reliability of this detector combination.
@@ 5 Conclusions
In this study, we introduced a novel approach for assessing the energy spectrum of extensive air showers (EAS) using a multi-hybrid detector station. By combining measurements from a scintillator surface detector (SSD), a water-Cherenkov detector (WCD), and resistive plate chambers (RPCs), we demonstrated the feasibility of disentangling the electromagnetic and muonic components of air showers at ground level.
@@
Overall, this work demonstrates that a single multi-hybrid station, when embedded within a sparse surface array, can provide valuable insights into the energy spectrum of shower components. These findings open new avenues for future experimental efforts, particularly in the context of hybrid cosmic ray observatories. Further studies with real data, along with refinements in the reconstruction algorithms, will be crucial in validating this methodology under actual observational conditions.
@@ Acknowledgements
We would like to thank Kevin Almeida Cheminant, Carola Dobrigkeit, Mário Pimenta and Jakub Vícha for carefully reading the manuscript and providing useful comments. This work has been funded by Fundação para a Ciência e Tecnologia, Portugal, under project https://doi.org/10.54499/2024.06879.CERN.

### v2

# Sensitivity of Hybrid Particle Detectors to the Energy Spectra of Air Shower Components

@@ 2.1 Simulation Framework
In this work, the extensive air showers were simulated using CORSIKA v7.7410 [21], and the detector response is obtained using a stand-alone program based on the Geant4 toolkit (v10.5.1) [22, 23]. This standalone simulation was qualitatively validated against the Auger simulation framework [24] for the three detector responses (SSD, WCD, and RPC), with the detector signals found to be in reasonable agreement. We used a reference configuration for the initial set of simulations, along with additional datasets to assess the robustness of the method employed here with respect to the reconstruction of key shower parameters, such as energy, zenith angle, and core position.
The reference configuration consists of CORSIKA proton-induced air showers with a fixed primary energy of E eV and a zenith angle of . A thinning algorithm with was applied and hadronic interactions were handled by FLUKA and EPOS-LHC at low and high energies, respectively. In each event, the shower core is injected at a fixed position relative to a specific detector station. This setup serves as a control configuration for our analysis.
To evaluate the dependence of the reconstruction on specific variables, we generated dedicated simulation sets where one parameter was varied at a time. For energy resolution studies, 5000 showers were simulated with energies ranging from eV to eV. For angular resolution, 2000 showers were generated at a zenith angle of . The shower core position was randomly distributed within an area of to investigate the impact of core position reconstruction.
@@ 2.2 Method to evaluate the sensitivity to the particle energy spectrum
In this work, we take the energy spectra of the electromagnetic and muonic components at ground level—obtained from CORSIKA simulations—and modify them while preserving the functional forms of their shapes. This approach is based on the assumption that differences in the energy spectra may arise from the limited ability of hadronic interaction models to accurately describe the underlying physics of shower development, rather than from the fundamental mechanisms of the shower itself. By doing so, we are able to investigate how variations in the particle energy distribution influence the detector response. In particular, we adjust the slopes of the low- and high-energy power-law tails while keeping the total particle multiplicity fixed.
The original particle energy distribution is obtained by processing CORSIKA-generated air showers down to the detector level using the aforementioned Geant4-based framework. Figure 3 illustrates examples of these distributions for the electromagnetic and muonic shower components.
To modify the shape of the distributions, we apply an energy-dependent weight in a selected energy range. The weighting function is defined as
where is the particle energy, and and define the lower and upper bounds of the energy window over which the modification is applied. The parameter controls the total weight increase (or decrease) across this window. This function modifies the spectrum’s tail in a controlled manner. Specifically, it ensures that at , and at , resulting in a smooth transition of the weights from 1 to within a selected energy range.
The choice of a monotonically increasing function for weighting the original spectrum is deliberate. Applying a flat increase within a limited region would introduce abrupt spikes at the transition between the modified and unmodified regions. This functional form ensures a smooth transition between the modified and unmodified regions, preserving the spectrum’s overall shape and features. After applying the weights, the full distribution is renormalized to conserve the total number of particles.
@@
To enable quantitative comparison across different detector types and avoid unit inconsistencies, the detector signals in the difference plots are normalized to their respective mean values from the original distributions. This yields dimensionless variables and , which represent the relative change in detector response and allow for direct comparison of sensitivity across different detector technologies.
@@
While and are not direct measurements of the spectrum itself, they provide a useful way to compare experimental data with simulations incorporating specific spectral modifications . By computing these variables from data and comparing them to those from simulated scenarios, we gain insight into the spectral changes that best reproduce the observed detector signals. In this way, and serve as effective intermediaries between the detector observables and the underlying energy spectra of the shower components, offering a practical method to interpret the data in terms of deviations from the reference simulation.
@@ 3.2 Low-energy tail of the muonic component
Following the approach used for modifying the high-energy tail of the electromagnetic particle energy distribution (Section 3.1), we considered a range of modifications with magnitudes from to . For each case, the estimators and were calculated for both SSD-WCD and WCD-RPC difference plots. The results are presented in Figure 10. Note that in this and all other radial plots shown in the article, the values of always lie within the principal angular range indicated by the plot axes.
In this case, the WCD-RPC difference plots demonstrate a much clearer separation of values across different modification intensities when compared with the SSD-WCD difference plots. This effect is due to the RPC detector exhibiting significantly higher sensitivity to changes in the muonic component compared to the WCD, while the SSD remains largely insensitive to muonic variations. This sensitivity difference comes from the fact that at this distance from the shower core, the electromagnetic and muonic contributions to the WCD signal remain comparable, whereas the RPC signal is clearly dominated by muons, as shown in Figure 2.
Alongside this, for the SSD-WCD difference plot, the values of are less consistent, particularly at lower values of . This is likely due to reduced statistical precision, a consequence of positioning the station farther from the shower core. Nevertheless, for the WCD-RPC difference plots, the values of remain consistent across all modification intensities, emphasizing the reliability of this detector combination.
@@ 5 Conclusions
In this study, we introduced a novel approach for assessing the energy spectrum of extensive air showers (EAS) components using a multi-hybrid detector station. By combining measurements from a scintillator surface detector (SSD), a water-Cherenkov detector (WCD), and resistive plate chambers (RPCs), we demonstrated the feasibility of disentangling the electromagnetic and muonic components of air showers at ground level.
@@
Overall, this work demonstrates that a single multi-hybrid station, when embedded within a sparse surface array, can provide valuable insights into the energy spectrum of shower components. These findings open new avenues for future experimental efforts, particularly in the context of hybrid cosmic ray observatories. For instance, within the Pierre Auger Observatory Infill array, the Underground Muon Detector (UMD) [29] could also contribute to a multi-hybrid configuration through its complementary sensitivity to muons. Its buried location suppresses the electromagnetic component more effectively and selects, on average, higher-energy muons than the water Cherenkov detectors, thus providing additional information. A dedicated study would, however, be required to properly assess its contribution, which is beyond the scope of the present work. Further studies with real data, along with refinements in the reconstruction algorithms, will be crucial in validating this methodology under actual observational conditions.
@@ Acknowledgements
We would like to thank Kevin Almeida Cheminant, Carola Dobrigkeit, Mário Pimenta and Jakub Vícha for carefully reading the manuscript and providing useful comments. We would also like to thank the Pierre Auger Collaboration for the helpful discussions. This work has been funded by Fundação para a Ciência e Tecnologia, Portugal, under project https://doi.org/10.54499/2024.06879.CERN. M.F. acknowledges financial support from FCT through grant 2025.00691.BD.